# RouteKG: A knowledge graph-based framework for route prediction on road networks

### v2

@@ 3.2 Knowledge Graph
4 Methodology
4.1 RouteKG Framework Overview
@@
Next, the Knowledge Graph Module is the core component of the proposed model, it takes the observed routes $\mathcal{X}^{o}$ and road network $\mathbf{G}$ to predict future routes. It constructs a knowledge graph on the road network $\mathbf{G}$, learns spatial relations $\mathcal{R}$, and predicts future routes, converting $\mathcal{X}^{o}$ to future route probabilities $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}})$ via $\left(\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathcal{R}\right)=\mathcal{M}_{kg}\left(\mathcal{X}^{o},\mathbf{G},\mathbf{D};\Theta_{kg}\right)$, where $\Theta_{kg}$ are the module’s parameters, and $\mathcal{R}$ represents the learned spatial relations.
With the future route probabilities $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}})$ in hand, the Route Generation Module employs an $n$-ary tree algorithm to generate potential future routes, yielding the top-$K$ preliminary route predictions. This step is captured by $\left\{\widetilde{\mathcal{X}^{f}_{k}}\right\}_{k=1}^{K}=\mathcal{M}_{g}\left(\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathbf{G},\mathbf{A}\right)$, where $\widetilde{\mathcal{X}^{f}_{k}}$ denotes the $k$-th generated future route.
When predicting future routes, predicted road links at different time-step are not independent but related. Thus, a Rank Refinement Module is utilized to collectively learn and assess the predicted route. It takes the initial top-$K$ predictions, $\left\{\widetilde{\mathcal{X}^{f}_{k}}\right\}_{k=1}^{K}$, and refines them using the spatial relations, $\mathcal{R}$. This refinement is achieved by the mapping $\left\{\widehat{\mathcal{X}^{f}_{k}}\right\}_{k=1}^{K}=\mathcal{M}_{r}\left(\left\{\widetilde{\mathcal{X}^{f}_{k}}\right\}_{k=1}^{K},\mathcal{R};\Theta_{r}\right)$, where $\Theta_{r}$ are the module’s parameters. This stage ensures that the final route predictions are accurate by considering the sequence of routes and spatial relations.
@@
4.2 Data Preprocessing Module
@@
Routes $\mathcal{X}$ are obtained by map-matching GPS trajectories $\mathcal{T}$ to the road network $\mathbf{G}$ (Yang & Gidofalvi 2018). These routes are then divided into observed routes $\mathcal{X}^{o}$ and future routes $\mathcal{X}^{f}$. Considering the importance of direction information in navigation (Chrastil & Warren 2015), we discretize continuous directions into $N_{d}$ classes to form $\mathcal{X}_{d}$ and $\mathbf{D}$. Figure 2 provides an example based on $N_{d}=8$. It has been shown that 8 directions are adequate in uniquely mapping most link-to-link movements and can enhance route prediction performance (Liang & Zhao 2021). This discretization allows for the convenient computation and assignment of inter- and intra-edge direction labels. It is worth noting that the two-way roads are given only one direction label for simplicity.
@@
To preserve the road network structure information, we construct the node adjacency edges (NAE) matrix, denoted as $\mathbf{A}$, can be derived directly from the road network $\mathbf{G}$. To build $\mathbf{A}$, we pad the edges adjacent to each node to an uniform length, thereby creating a matrix of dimensions $\mathbb{R}^{|\mathbf{V}|\times N_{A}}$. In this context, $|\mathbf{V}|$ indicates the total number of nodes in the road network, while $N_{A}$ represents the maximum number of adjacent edges to any node. This padding approach enables batch training combined with smart masking techniques.
4.3 Knowledge Graph Module
@@
After data preprocessing, we designed a Knowledge Graph Module that adapts the KG to the road network, which learns the complex spatial relationships between road links and therefore more accurately estimate the probability of each link as part of the future route $x^{f}$, given an observed route $x^{o}$. Formally, given a road network $\mathbf{G}$ and an observed route $x^{o}\in\mathcal{X}^{o}$, the module outputs $\Gamma^{\prime}$ probability distributions $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}})=\left\{\mathrm{Pr}\left(\widetilde{x^{f,\gamma}}\right)\right\}_{\gamma=1}^{\Gamma^{\prime}}$. Each distribution indicates the probability of a link being part of future routes, with the $\gamma$-th distribution indicating the likelihood of each road link being the $\gamma$-th link in those future routes, where $\gamma=1,2,\dots,\Gamma^{\prime}$.
@@
Intuitively, a driver’s route choice is based on their intended goal. Therefore, using KGC for route prediction aligns with the logic behind drivers’ route selections. However, most existing KGs are designed for search engines (Xiong et al. 2017) and text-based Question Answering (Huang et al. 2019), making them unsuitable for direct application to road networks. Therefore, we need to construct a KG tailored to the characteristics of road networks, redefine the KGC problem in this context, and use learned spatial relations for more accurate route prediction. These tasks are encompassed in three submodules we’ve designed: Knowledge Graph Construction, Knowledge Graph Representation Learning, and Future Route Prediction through KGC. We will detail these in the following subsections.
4.3.1 Knowledge Graph Construction
@@
To design a KG $\mathcal{G}$ tailored for road networks and route prediction, we first need to select the crucial spatial and structural features in road networks. The desired KG should preserve the spatial relations amongst the identified entities while maintaining its applicability and generalizability across fine-grained scenarios on the road networks. In alignment with this objective, the selection focuses solely on those entities and relations that pervade all road networks and routing contexts. A detailed explanation of the entity and relation selection processes is provided below.
@@
When constructing the KG $\mathcal{G}$ for road networks and routes, the initial key step is to identify entities $\mathcal{E}$. In the context of a road network, the predominant entity is the link. Every link $e$ is characterized by their unique identifiers and associated attributes such as length or connectivity. Selecting links as the sole entities reflects their intrinsic importance within the road network. It ensures the broad generalizability of the resulting KG across various routing contexts and scenarios, contributing to the applicability of the proposed approach.
@@
The four major spatial relation types are summarized in Table 1. By comprehensively capturing these four types of relations over the identified entities, the KG possesses a rich and nuanced representation of the road network, which could facilitate various advanced routing tasks.
Table 1: Major spatial relation types and corresponding notations and data sources of the constructed KG. A major relation type may contain multiple relations. For example, the relation type “DirectionTo” contains $N_{d}$ directions, indicating a total of $N_{d}$ direction relations.
@@
4.3.2 Knowledge Graph Representation Learning
Spatial relations between entities (i.e. links) on a road network should be route-agnostic. This means that these relations should be independent of specific routes and instead solely reflect the spatial attributes of the road network itself. These relations also need to be encoded efficiently to support training and inference. One common approach is to employ KG embedding techniques, which aims to find embedding functions $\mathcal{M}_{\mathcal{E}}$, $\mathcal{M}_{\mathcal{R}}$ that maps each entity and each relation in to an feature vector. The embedding function $\mathcal{M}_{\mathcal{E}}(\cdot)$ and $\mathcal{M}_{\mathcal{E}}(\cdot)$ should preserve the inherent property of $\mathcal{G}$. However, road networks exhibit complex relations, as illustrated in Figure 3, involving many-to-many relations. To address this complexity, we modify and adapt the translation distance model TransH (Bordes et al. 2013) to $\mathcal{G}$ in our study. This enables us to effectively learn the vector representations of both entities and relations in the knowledge graph $\mathcal{G}$.
@@ 4.3.2 Knowledge Graph Representation Learning
To enable KG representation learning, we first need to construct the sets of positive triplets $\Delta$ and negative triplets $\Delta^{\prime}$ for each relation type. To facilitate batch representation learning and ensure comprehensive learning of all entities and relations, we employ a Random Sampling-based method for KG representation learning. The subsequent paragraphs will offer detailed description of the positive and negative triplet sets’ construction processes for each relation type in $\mathcal{G}$.
@@
$\mathcal{L}_{rep}=\sum_{\Delta,\Delta^{\prime}\in\left\{(\Delta_{\mathcal{R}^{\cdot}},\Delta_{\mathcal{R}^{\cdot}}^{\prime})\right\}}\sum_{(h,r,t)\in\Delta}\sum_{(h^{\prime},r^{\prime},t^{\prime})\in\Delta^{\prime}}\Biggl{[}$  $\left(\mathbf{h}-\left(\mathbf{p}^{r}\right)^{\top}\mathbf{h}\mathbf{p}^{r}\right)+\mathbf{r}-\left(\mathbf{t}-\left(\mathbf{p}^{r}\right)^{\top}\mathbf{t}\mathbf{p}^{r}\right)\text{ }\right\|_{\ell_{1}}+\psi-$ (4)
 $\left(\mathbf{h}^{\prime}-\left(\mathbf{p}^{r^{\prime}}\right)^{\top}\mathbf{h}^{\prime}\mathbf{p}^{r^{\prime}}\right)+\mathbf{r}^{\prime}-\left(\mathbf{t}^{\prime}-\left(\mathbf{p}^{r^{\prime}}\right)^{\top}\mathbf{t}^{\prime}\mathbf{p}^{r^{\prime}}\right)\text{ }\right\|_{\ell_{1}}\text{ }\Biggr{]}_{+},$
@@
4.3.3 Future Route Prediction through Knowledge Graph Completion
@@
For illustration, we consider the $i$-th observed route $x^{o}_{i}\in\mathcal{X}^{o}$ and the corresponding direction $x^{o,d}_{i}\in\mathcal{X}^{o,d}$, drawn from the set of observed routes $\mathcal{X}^{o}$ and directions $\mathcal{X}^{o,d}$, respectively. Here, $x^{o}_{i}=\{e^{j}_{i}\}^{\Gamma}_{j=1}$ and $x^{o,d}_{i}=\{e^{d,j}_{i}\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$. Initially, we extract the embeddings of all elements of $x^{o}$ and $x^{o,d}$ by multiplying their respective one-hot vectors with the corresponding trainable embedding matrices: $\mathbf{W}_{\mathcal{E}}\in\mathbb{R}^{|\mathcal{E}|\times\delta_{\mathcal{E}}}$ and $\mathbf{W}_{\mathcal{R}^{d}}\in\mathbb{R}^{|\mathcal{R}^{d}|\times\delta_{\mathcal{R}^{d}}}$. The resulting embeddings for $x^{o}_{i}$ and $x^{o,d}_{i}$ are denoted as $\mathbf{x}^{o}_{i}=\{\mathbf{e}^{j}_{i}\|\}_{j=1}^{\Gamma}$ and $\mathbf{x}^{o,d}_{i}=\{\mathbf{e}^{d,j}_{i}\|\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$, with $\mathbf{e}_{i}^{\cdot}\in\mathbb{R}^{\delta_{\mathcal{E}}}$, $\mathbf{e}^{d,\cdot}_{i}\in\mathbb{R}^{\delta_{\mathcal{R}^{d}}}$. To predict the direction of the vehicle’s future routes, we utilize a Multi-layer Perceptron (MLP) (Popescu et al. 2009) to encode $\mathbf{x}^{o}_{i}$ and $\mathbf{x}^{o,d}_{i}$:
$\widehat{r^{d}_{i}}=\operatorname*{argmax}\left(\textsc{MLP}_{d}\left(\mathbf{x}^{o}_{i}\text{ }\|\text{ }\mathbf{x}^{o,d}_{i}\right)\right),$ (5)
@@
$\mathrm{Pr}\left(\widetilde{x^{f,\gamma}_{i}}\right)=\mathrm{Softmax}\left(\mathbf{e}_{\perp}\cdot\left[\left(\mathbf{e}_{i,\perp}^{\Gamma}+\mathbf{r}^{d}_{i}\right)\odot\mathbf{r}^{a,\gamma}_{i}\right]^{\top}\right),$ (8)
where $\mathrm{Pr}\left(\widetilde{x^{f,\gamma}_{i}}\right)\in\mathbb{R}^{|\mathcal{E}|}$ is the predicted probability distribution which indicates the likelihood of each link being the $\gamma$-th link of the $i$-th future route. We can set $\gamma$ from 1 to $\Gamma^{\prime}$ and recursively use Eq. (6) to obtain $\Gamma^{\prime}$ probability distributions $\left\{\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})\right\}_{\gamma=1}^{\Gamma^{\prime}}$ representing the estimated future route probabilities, which is the final output of the module.
@@
$\mathrm{Pr}\left(\widetilde{x^{f,\gamma}_{i}}\right)=\mathrm{Softmax}\left(\mathbf{e}_{\perp}\cdot\left[\left(\mathbf{e}_{i,\perp}^{\Gamma}+\mathbf{e}_{i,\perp}^{\Gamma+\Gamma^{\prime}}+\mathbf{r}^{d}_{i}\right)\odot\mathbf{r}^{a,\gamma}_{i}\right]^{\top}\right),$ (9)
@@
$\mathcal{L}_{pred}=-\sum_{i=1}^{|\mathcal{X}|}\sum_{\gamma=1}^{\Gamma^{\prime}}\log\mathrm{Pr}\left(\widetilde{x^{f,\gamma}_{i}}\right)\left[x_{i}^{f,\gamma}\right],$ (10)

### v3

@@ III-B Knowledge Graph
IV Methodology
IV-A RouteKG Framework Overview
@@
Next, the Knowledge Graph Module is the core component of the proposed model, it takes the observed routes $\mathcal{X}^{o}$ and road network $\mathbf{G}$ to predict future routes. It constructs a knowledge graph on the road network $\mathbf{G}$, learns spatial relations $\mathcal{R}$, and predicts future routes, converting $\mathcal{X}^{o}$ to future route probabilities $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}})$ via $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathcal{R}=\mathcal{M}_{kg}\left(\mathcal{X}^{o},\mathbf{G},\mathbf{D};\Theta_{kg}\right)$, where $\Theta_{kg}$ are the module’s parameters, and $\mathcal{R}$ represents the learned spatial relations.
With the future route probabilities $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}})$ in hand, the Route Generation Module employs an $n$-ary tree algorithm to generate potential future routes, yielding the top-$K$ preliminary route predictions. This step is captured by $\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}=\mathcal{M}_{g}(\mathrm{Pr}(\widetilde{\mathcal{X}^{f}}),\mathbf{G},\mathbf{A})$, where $\widetilde{\mathcal{X}^{f}_{k}}$ denotes the $k$-th generated future route.
When predicting future routes, predicted road links at different time steps are not independent but related. Thus, a Rank Refinement Module is utilized to collectively learn and assess the predicted route. It takes the initial top-$K$ predictions, $\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}$, and refines them using the spatial relations, $\mathcal{R}$. This refinement is achieved by the mapping $\{\widehat{\mathcal{X}^{f}_{k}}\}_{k=1}^{K}=\mathcal{M}_{r}(\{\widetilde{\mathcal{X}^{f}_{k}}\}_{k=1}^{K},\mathcal{R};\Theta_{r})$, where $\Theta_{r}$ are the module’s parameters. This stage ensures that the final route predictions are accurate by considering the sequence of routes and spatial relations.
@@
IV-B Data Preprocessing Module
@@
Routes $\mathcal{X}$ are obtained by map-matching GPS trajectories $\mathcal{T}$ to the road network $\mathbf{G}$ [72]. These routes are then divided into observed routes $\mathcal{X}^{o}$ and future routes $\mathcal{X}^{f}$. Considering the importance of direction information in navigation [19], we discretize continuous directions into $N_{d}$ classes to form $\mathcal{X}_{d}$ and $\mathbf{D}$. Figure 2 provides an example based on $N_{d}=8$. It has been shown that 8 directions are adequate in uniquely mapping most link-to-link movements and can enhance route prediction performance [14]. This discretization allows for the convenient computation and assignment of inter- and intra-edge direction labels. It is worth noting that the two-way roads are given only one direction label for simplicity.
@@
To preserve the road network structure information, we construct the node adjacency edges (NAE) matrix, denoted as $\mathbf{A}$, which can be derived directly from the road network $\mathbf{G}$. To build $\mathbf{A}$, we pad the edges adjacent to each node to a uniform length, thereby creating a matrix of dimensions $\mathbb{R}^{|\mathbf{V}|\times N_{A}}$. In this context, $|\mathbf{V}|$ indicates the total number of nodes in the road network, while $N_{A}$ represents the maximum number of adjacent edges to any node. This padding approach enables batch training combined with smart masking techniques.
IV-C Knowledge Graph Module
@@
After data preprocessing, we designed a Knowledge Graph Module that adapts the KG to the road network, which learns the complex spatial relationships between road links and therefore more accurately estimates the probability of each link as part of the future route $x^{f}$, given an observed route $x^{o}$. Formally, given a road network $\mathbf{G}$ and an observed route $x^{o}\in\mathcal{X}^{o}$, the module outputs $\Gamma^{\prime}$ probability distributions $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}})=\left\{\mathrm{Pr}(\widetilde{x^{f,\gamma}})\right\}_{\gamma=1}^{\Gamma^{\prime}}$. Each distribution indicates the probability of a link being part of future routes, with the $\gamma$-th distribution indicating the likelihood of each road link being the $\gamma$-th link in those future routes, where $\gamma=1,2,\dots,\Gamma^{\prime}$.
@@
Intuitively, a driver’s route choice is based on their intended goal. Therefore, using KGC for route prediction aligns with the logic behind drivers’ route selections. However, most existing KGs are designed for search engines [73] and text-based Question Answering [74], making them unsuitable for direct application to road networks. Therefore, we need to construct a KG tailored to the characteristics of road networks, redefine the KGC problem in this context, and use learned spatial relations for more accurate route prediction. These tasks are encompassed in three submodules we’ve designed: Knowledge Graph Construction, Knowledge Graph Representation Learning, and Future Route Prediction through KGC. We will detail these in the following subsections.
IV-C1 Knowledge Graph Construction
@@
To design a KG $\mathcal{G}$ tailored for road networks and route prediction, we first need to select the crucial spatial and structural features in road networks. The desired KG should preserve the spatial relations amongst the identified entities while maintaining its applicability and generalizability across fine-grained scenarios on road networks. In alignment with this objective, the selection focuses solely on those entities and relations that pervade all road networks and routing contexts. A detailed explanation of the entity and relation selection processes is provided below.
@@
When constructing the KG $\mathcal{G}$ for road networks and routes, the initial key step is to identify entities $\mathcal{E}$. In the context of a road network, the predominant entity is the link. Every link $e$ is characterized by its unique identifiers and associated attributes such as length or connectivity. Selecting links as the sole entities reflects their intrinsic importance within the road network. It ensures the broad generalizability of the resulting KG across various routing contexts and scenarios, contributing to the applicability of the proposed approach.
@@
Table I summarizes the four major spatial relation types. By comprehensively capturing these four types of relations over the identified entities, the KG possesses a rich and nuanced representation of the road network, which could facilitate various advanced routing tasks.
@@
TABLE I: Major spatial relation types and corresponding notations and data sources of the constructed KG. A major relation type may contain multiple relations. For example, the relation type “DirectionTo” contains $N_{d}$ directions, indicating a total of $N_{d}$ direction relations.
IV-C2 Knowledge Graph Representation Learning
Spatial relations between entities (i.e. links) on a road network should be route-agnostic. This means that these relations should be independent of specific routes and instead solely reflect the spatial attributes of the road network itself. These relations also need to be encoded efficiently to support training and inference. One common approach is to employ KG embedding techniques, which aims to find embedding functions $\mathcal{M}_{\mathcal{E}}$, $\mathcal{M}_{\mathcal{R}}$ that maps each entity and each relation in to an feature vector. The embedding function $\mathcal{M}_{\mathcal{E}}(\cdot)$ and $\mathcal{M}_{\mathcal{E}}(\cdot)$ should preserve the inherent property of $\mathcal{G}$. However, road networks exhibit complex relations, as illustrated in Figure 3, involving many-to-many relations. To address this complexity, we modify and adapt the translation distance model TransH [52] to $\mathcal{G}$ in our study. This enables us to effectively learn the vector representations of both entities and relations in the knowledge graph $\mathcal{G}$.
@@ IV-C2 Knowledge Graph Representation Learning
To enable KG representation learning, we first need to construct the sets of positive triplets $\Delta$ and negative triplets $\Delta^{\prime}$ for each relation type. To facilitate batch representation learning and ensure comprehensive learning of all entities and relations, we employ a Random Sampling-based method for KG representation learning. The subsequent paragraphs will offer detailed descriptions of the positive and negative triplet sets’ construction processes for each relation type in $\mathcal{G}$.
@@
$\mathcal{L}_{rep}=\sum_{\Delta,\Delta^{\prime}\in\left\{(\Delta_{\mathcal{R}^{\cdot}},\Delta_{\mathcal{R}^{\cdot}}^{\prime})\right\}}\sum_{(h,r,t)\in\Delta}\sum_{(h^{\prime},r^{\prime},t^{\prime})\in\Delta^{\prime}}$ (4)
[ $\left(\mathbf{h}-\left(\mathbf{p}^{r}\right)^{\top}\mathbf{h}\mathbf{p}^{r}\right)+\mathbf{r}-\left(\mathbf{t}-\left(\mathbf{p}^{r}\right)^{\top}\mathbf{t}\mathbf{p}^{r}\right)\text{ }\right\|_{\ell_{1}}+\psi-$
 $\left(\mathbf{h}^{\prime}-(\mathbf{p}^{r^{\prime}})^{\top}\mathbf{h}^{\prime}\mathbf{p}^{r^{\prime}}\right)+\mathbf{r}^{\prime}-\left(\mathbf{t}^{\prime}-(\mathbf{p}^{r^{\prime}})^{\top}\mathbf{t}^{\prime}\mathbf{p}^{r^{\prime}}\right)\text{ }\right\|_{\ell_{1}}\text{ }\Biggr{]}_{+},$
@@
IV-C3 Future Route Prediction through Knowledge Graph Completion
@@
For illustration, we consider the $i$-th observed route $x^{o}_{i}\in\mathcal{X}^{o}$ and the corresponding direction $x^{o,d}_{i}\in\mathcal{X}^{o,d}$, drawn from the set of observed routes $\mathcal{X}^{o}$ and directions $\mathcal{X}^{o,d}$, respectively. Here, $x^{o}_{i}=\{e^{j}_{i}\}^{\Gamma}_{j=1}$ and $x^{o,d}_{i}=\{e^{d,j}_{i}\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$. Initially, we extract the embeddings of all elements of $x^{o}$ and $x^{o,d}$ by multiplying their respective one-hot vectors with the corresponding trainable embedding matrices: $\mathbf{W}_{\mathcal{E}}\in\mathbb{R}^{|\mathcal{E}|\times\delta_{\mathcal{E}}}$ and $\mathbf{W}_{\mathcal{R}^{d}}\in\mathbb{R}^{|\mathcal{R}^{d}|\times\delta_{\mathcal{R}^{d}}}$. The resulting embeddings for $x^{o}_{i}$ and $x^{o,d}_{i}$ are denoted as $\mathbf{x}^{o}_{i}=\{\mathbf{e}^{j}_{i}\|\}_{j=1}^{\Gamma}$ and $\mathbf{x}^{o,d}_{i}=\{\mathbf{e}^{d,j}_{i}\|\}_{j=\Gamma+1}^{\Gamma+\Gamma^{\prime}}$, with $\mathbf{e}_{i}^{\cdot}\in\mathbb{R}^{\delta_{\mathcal{E}}}$, $\mathbf{e}^{d,\cdot}_{i}\in\mathbb{R}^{\delta_{\mathcal{R}^{d}}}$. To predict the direction of the vehicle’s future routes, we utilize a Multi-layer Perceptron (MLP) [75] to encode $\mathbf{x}^{o}_{i}$ and $\mathbf{x}^{o,d}_{i}$:
$\widehat{r^{d}_{i}}=\arg\max\left(\textsc{MLP}_{d}\left(\mathbf{x}^{o}_{i}\text{ }\|\text{ }\mathbf{x}^{o,d}_{i}\right)\right),$ (5)
@@
$\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})=\mathrm{Softmax}\left(\mathbf{e}_{\perp}\cdot\left[\left(\mathbf{e}_{i,\perp}^{\Gamma}+\mathbf{r}^{d}_{i}\right)\odot\mathbf{r}^{a,\gamma}_{i}\right]^{\top}\right),$ (8)
where $\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})\in\mathbb{R}^{|\mathcal{E}|}$ is the predicted probability distribution which indicates the likelihood of each link being the $\gamma$-th link of the $i$-th future route. We can set $\gamma$ from 1 to $\Gamma^{\prime}$ and recursively use Eq. (6) to obtain $\Gamma^{\prime}$ probability distributions $\{\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})\}_{\gamma=1}^{\Gamma^{\prime}}$ representing the estimated future route probabilities, which is the final output of the module.
@@
$\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})=\mathrm{Softmax}(\mathbf{e}_{\perp}\cdot\left[(\mathbf{e}_{i,\perp}^{\Gamma}+\mathbf{e}_{i,\perp}^{\Gamma+\Gamma^{\prime}}+\mathbf{r}^{d}_{i})\odot\mathbf{r}^{a,\gamma}_{i}\right]^{\top}),$ (9)
@@
$\mathcal{L}_{pred}=-\sum_{i=1}^{|\mathcal{X}|}\sum_{\gamma=1}^{\Gamma^{\prime}}\log\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})\left[x_{i}^{f,\gamma}\right],$ (10)
@@
IV-D Route Generation Module
Given predicted future route probabilities $\mathrm{Pr}(\widetilde{\mathcal{X}^{f}})$, the Route Generation Module generates multiple possible future routes from these probabilities. An $n$-ary tree-based algorithm, Spanning Route, is proposed to generate these future routes based on the predicted probabilities. This algorithm is visualized in Figure 5 through a simplified case where $n=2$ and only $\Gamma=3$ predicted future links are illustrated. For each tree node, we designate four attributes: name, parent, end_node, and pred. The name corresponds to the identification of the leaf, while the parent points to the predecessor of the current leaf. The attribute end_node signifies the terminal node of the current predicted link, and pred is the present predictions.